# Interacting binaries on the Main Sequence as in-situ tracers of mass transfer efficiency and stability

### v1

@@ 4.4 Wind mass loss
The wind mass loss rate in the detailed binary evolution models are set as in I. Brott et al. (2011). The mass loss rate on the Main Sequence is low compared to that of evolved systems and also decreases with decreasing metallicity (J. S. Vink et al., 2001; M. R. Mokiem et al., 2007; M. Renzo et al., 2017; R. Björklund et al., 2021; C. Hawcroft et al., 2021; S. A. Brands et al., 2022; A. C. Gormaz-Matamala et al., 2023; J. Krticka et al., 2025). The donors in the binary models typically lose $\leq 1$$M_{\odot}$ via stellar winds before the thermal timescale mass transfer phase initiates on the Main Sequence (I. Brott et al., 2011; K. Sen et al., 2022). Uncertainties in the wind mass-loss rates can also impact these predicted core masses, but this effect is small for initial masses below $\simeq 30M_{\odot}$: at solar metallicity, the corresponding systematic error on Helium core masses is $\lesssim 0.4M_{\odot}$ (M. Renzo et al., 2017). The amount of mass lost via winds is lower for SMC metallicity. The mass transfer rate is orders of magnitude higher than the wind mass loss rate, such that our conclusions on mass transfer efficiency are not significantly affected by ignoring the mass lost by the donor and accretor via winds during the mass transfer phases.
4.5 Observational sample
@@
The current sample of massive Algol binaries is highly heterogeneous. The component masses for the Galactic Algols were taken from individual studies, and the data analysis techniques vary from one study to another. Moreover, most of the observational data were taken several decades ago, without recent spectroscopic follow-ups. The mass estimates are often without error bars. Only four out of the seven massive Algols in the LMC have estimates on surface abundances and rotational velocities of the individual components. Of the eight massive Algols in the LMC and the Milky Way that have surface rotation measurements, only two show donors that are rotating synchronously with the orbit (K. Sen et al., 2022), while our current binary evolution models predict that all Roche-lobe filling donors should be synchronised to the orbit.
@@
None of the SMC systems has surface abundance or rotational velocity measurements either. The effective temperature measurements of the primary components of the massive Algols in the SMC were based on spectral type classification and the effective temperature of the binary companion was derived from the I-band flux ratios (T. J. Harries et al., 2003; R. W. Hilditch et al., 2005), which lie in the Rayleigh-Jeans tail and are not the most sensitive probes of the effective temperatures of the O- and B-type stars. Therefore, modern spectroscopic surveys of massive Algols binaries are essential to obtain accurate information on most of the existing literature, beyond the component masses and the orbital periods. Moreover, a larger number of observational constraints, such as rotational velocities and surface abundances, can be crucial to understand the extent of envelope stripping and the mode of mass transfer (disk or no disk) occurring in these massive Algols. Like the OGLE survey, the ASAS-SN survey (C. S. Kochanek et al., 2017) can be an excellent dataset to mine eclipsing semi-detached binaries for follow-up homogeneous spectroscopic monitoring campaigns in the Milky Way and the Magellanic Clouds.
5 Conclusion
@@
We investigated the efficiency, angular momentum loss, and stability of mass transfer in massive interacting binaries on the Main Sequence (Algols) across the Milky Way, Large Magellanic Cloud (LMC), and Small Magellanic Cloud (SMC). Using an analytical framework that connects present-day binary properties (donor mass, accretor mass, and orbital period) to their initial parameters, and combining these with physical constraints, we derived the ranges of mass transfer efficiency and angular momentum loss consistent with the observed populations. Our main conclusions are:
@@
  • The current set of massive Algols requires that mass transfer remain stable down to initial mass ratios of $\sim$0.6, lower than typically assumed in the literature (S. E. de Mink et al., 2013; H. Ge et al., 2020) for mass transfer on the Main Sequence. The least massive Algols ($M_{\rm d,i}<20\,M_{\odot}$) provide the most substantial evidence for these relaxed stability criteria.
@@
Multi-epoch spectroscopic surveys of massive stars and binaries (e.g. the OWN survey, R. H. Barbá et al., 2017; the IACOB project, S. Simón-Díaz & A. Herrero, 2014; the VFTS survey, C. J. Evans et al., 2011; the BLOeM survey, T. Shenar et al., 2024), especially with constraints from rotational velocities and surface abundances, are crucial for further disentangling the degeneracy between efficiency and stability, and for anchoring population synthesis predictions of stripped-envelope supernovae and gravitational-wave sources. Simultaneous improvements in multi-dimensional modelling of the thermal timescale mass transfer (D. V. Bisikalo et al., 1998; L. Dessart et al., 2003; W. Lu et al., 2023; J. Cehula & O. Pejcha, 2023; T. Ryu et al., 2025; P. Scherbak et al., 2025a, b) and improved orbital modelling in 1D codes (K. A. Rocha et al., 2025; A. Parkosidis et al., 2025) will help advance the field, enabling more accurate modelling of the early stages of mass transfer.
M.R. acknowledges support from NASA (ATP: 80NSSC24K0932). JIV acknowledges support from the European Research Council for the ERC Advanced Grant 101054731. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). Part of this work was initiated at the ”Stable mass transfer 2.0” workshop held at the Center for Computational Astrophysics of the Flatiron Institute, which is supported by the Simons Foundation. We thank Pablo Marchant and Tom Maccarone for their comments on the manuscript.
@@ Appendix A Additional plots and tables
Table 3: Stellar and binary parameters used in this work (see Sect. 2) taken from detailed binary evolution models for the LMC and SMC.

### v2

@@ IV.4 Wind mass loss
The wind mass loss rate in the detailed binary evolution models are set as in Brott et al. (2011). The mass loss rate on the Main Sequence is low compared to that of evolved systems and also decreases with decreasing metallicity (Vink et al., 2001; Mokiem et al., 2007; Renzo et al., 2017; Björklund et al., 2021; Hawcroft et al., 2021; Brands et al., 2022; Gormaz-Matamala et al., 2023; Krticka et al., 2025). The donors in the binary models typically lose $\leq 1$$M_{\odot}$ via stellar winds before the thermal timescale mass transfer phase initiates on the Main Sequence (Brott et al., 2011; Sen et al., 2022). Uncertainties in the wind mass-loss rates can also impact these predicted core masses, but this effect is small for initial masses below $\simeq 30M_{\odot}$: at solar metallicity, the corresponding systematic error on Helium core masses is $\lesssim 0.4M_{\odot}$ (Renzo et al., 2017). The amount of mass lost via winds is lower for SMC metallicity. The mass transfer rate is orders of magnitude higher than the wind mass loss rate, such that our conclusions on mass transfer efficiency are not significantly affected by ignoring the mass lost by the donor and accretor via winds during the mass transfer phases.
IV.5 Observational sample
@@
The current sample of massive Algol binaries is highly heterogeneous. The component masses for the Galactic Algols were taken from individual studies, and the data analysis techniques vary from one study to another. Moreover, most of the observational data were taken several decades ago, without recent spectroscopic follow-ups. The mass estimates are often without error bars. Only four out of the seven massive Algols in the LMC have estimates on surface abundances and rotational velocities of the individual components. Of the eight massive Algols in the LMC and the Milky Way that have surface rotation measurements, only two show donors that are rotating synchronously with the orbit (Sen et al., 2022), while our current binary evolution models predict that all Roche-lobe filling donors should be synchronised to the orbit.
@@
None of the SMC systems has surface abundance or rotational velocity measurements either. The effective temperature measurements of the primary components of the massive Algols in the SMC were based on spectral type classification and the effective temperature of the binary companion was derived from the I-band flux ratios (Harries et al., 2003; Hilditch et al., 2005), which lie in the Rayleigh-Jeans tail and are not the most sensitive probes of the effective temperatures of the O- and B-type stars. Therefore, modern spectroscopic surveys of massive Algols binaries are essential to obtain accurate information on most of the existing literature, beyond the component masses and the orbital periods. Moreover, a larger number of observational constraints, such as rotational velocities and surface abundances, can be crucial to understand the extent of envelope stripping and the mode of mass transfer (disk or no disk) occurring in these massive Algols. Like the OGLE survey, the ASAS-SN survey (Kochanek et al., 2017) can be an excellent dataset to mine eclipsing semi-detached binaries for follow-up homogeneous spectroscopic monitoring campaigns in the Milky Way and the Magellanic Clouds.
V Conclusion
@@
We investigated the efficiency, angular momentum loss, and stability of mass transfer in massive interacting binaries on the Main Sequence (Algols) across the Milky Way, Large Magellanic Cloud, and Small Magellanic Cloud. Using an analytical framework that connects present-day binary properties (donor mass, accretor mass, and orbital period) to their initial parameters, and combining these with physical constraints, we derived the ranges of mass transfer efficiency and angular momentum loss consistent with the observed populations. Our main conclusions are:
@@
  • The current set of massive Algols requires that mass transfer remain stable down to initial mass ratios of $\sim$0.6, lower than typically assumed in the literature (de Mink et al., 2013; Ge et al., 2020) as well as in our detailed binary evolution models (Wang et al., 2020; Sen et al., 2022). The least massive Algols ($M_{\rm d,i}<20\,M_{\odot}$) provide the most substantial evidence for these relaxed stability criteria.
  • Our analysis relies on the assumption (Sect. II.2 and Sect. IV.2) that the fast thermal-timescale initial mass-transfer strips the donor down to its Zero Age Main Sequence convective core mass. Some massive Algols cannot be reproduced by the detailed models (e.g. right panel of Fig. 1), which assume rotationally limited accretion efficiency and efficient tides. This implies that the mass-transfer efficiency may not entirely be regulated by rotational processes. Observations of massive Algols in the earliest post-mass-transfer phase may require improving the physical picture of mass transfer in detailed stellar evolution models.
@@
Multi-epoch spectroscopic surveys of massive stars and binaries (e.g. the OWN survey, Barbá et al., 2017; the IACOB project, Simón-Díaz and Herrero, 2014; the VFTS survey, Evans et al., 2011; the BLOeM survey, Shenar et al., 2024), especially with constraints from rotational velocities and surface abundances, are crucial for further disentangling the degeneracy between efficiency and stability, and for anchoring population synthesis predictions of stripped-envelope supernovae and gravitational-wave sources. Simultaneous improvements in multi-dimensional modelling of the thermal timescale mass transfer (Bisikalo et al., 1998; Dessart et al., 2003; Lu et al., 2023; Cehula and Pejcha, 2023; Ryu et al., 2025; Scherbak et al., 2025a, b) and improved orbital modelling in 1D codes (Rocha et al., 2025; Parkosidis et al., 2025) will help advance the field, enabling more accurate modelling of the early stages of mass transfer.
M.R. acknowledges support from NASA (ATP: 80NSSC24K0932). JIV acknowledges support from the European Research Council for the ERC Advanced Grant 101054731. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). Part of this work was initiated at the ”Stable mass transfer 2.0” workshop held at the Center for Computational Astrophysics of the Flatiron Institute, which is supported by the Simons Foundation. We thank Pablo Marchant, Tom Maccarone, Hugues Sana and the anonymous referee for helpful comments on the manuscript.
@@ Appendix A Additional plots and tables
Table 4: Stellar and binary parameters used in this work (see Sect. II) taken from detailed binary evolution models for the LMC and SMC.